OpenStudy (anonymous):

(4x^3 - 4x^2 - 9x + 9) ÷ (x + 3) help plz?

myininaya (myininaya):

---------------- x+3 | 4x^4-4x^2-9x+9 how many times does x go into 4x^4?

OpenStudy (anonymous):

uh? I think 4.

myininaya (myininaya):

what about the x part how many times does x go into x^4?

myininaya (myininaya):

\[\frac{x^4}{x}=\frac{x \cdot x \cdot x \cdot x}{x}\]

myininaya (myininaya):

does anything cancel?

OpenStudy (anonymous):

those x's?? im kinda confused.

OpenStudy (anonymous):

right side?

myininaya (myininaya):

yes those are x's

OpenStudy (anonymous):

@myininaya, you added an x when you copied the equation. It's \[4x^{3}\]not\[4x^{4}\]

OpenStudy (anonymous):

is -108 the remainder???

myininaya (myininaya):

oh okay how many times does x go into 4x^3? \[\frac{4x^3}{x}=\frac{4 \cdot x \cdot x \cdot x}{x}\] what cancels @j3rseykid1 ?

myininaya (myininaya):

There is also synthetic division if long division for this one

myininaya (myininaya):

Have you learned synthetic division?

OpenStudy (anonymous):

a little. im still learning it.

myininaya (myininaya):

Do you know how to set it up?

OpenStudy (anonymous):

yes

myininaya (myininaya):

Is that how you would set it up: -3 | 4 -4 -9 9 | ------------------ | We didn't need any zero placeholders because the exponents didn't skip any integers what i mean is we have 3,2,1,0 No integers skipped

OpenStudy (anonymous):

yes, i set it up like that.

myininaya (myininaya):

great do you know the next step?

OpenStudy (anonymous):

i think you would bring the 3 down and do

myininaya (myininaya):

what 3 there is no 3 there you mean 4? if so that is correct

OpenStudy (anonymous):

oh yes i ment 4..lol.

OpenStudy (anonymous):

Make sure you are using the correct equation! @j3rseykid1 said 4x^3, but @myininaya said 4x^4 in her first post. Don't solve for the wrong equation!

OpenStudy (anonymous):

I guess ima throw in the white flag on this one...#CLOSE

myininaya (myininaya):

yes you bring down the 4

myininaya (myininaya):

-3 | 4 -4 -9 9 | ------------------ 4 | Next step is?